OpenStudy (anonymous):

you have enough shrubs to cover an area of 100ft^2. What is the radius of the largest circular region you can plant with these shrubs? Round your answer to the nearest tenth of a foot.

zepdrix (zepdrix):



zepdrix (zepdrix):

\[\Large\rm A=\pi r^2\]\[\Large\rm 100=\pi r^2\]\[\Large\rm r=?\]

OpenStudy (anonymous):

50?